# Nonstabilizerness of a Boundary Time Crystal

### v1

@@ VI Conclusions
At the critical point of the measurement-induced phase transition, the two complexity measures reveal complementary aspects of the behavior of the system. The entanglement entropy exhibits a genuine phase transition with unraveling-dependent scaling, while magic displays a cusp singularity in its derivative while, importantly, remaining extensive in system size across both phases. This result suggests that the superradiant model hosts highly nonclassical resources, even in the absence of entanglement.
Figure 9: Histograms showing the probability density of the (a) magic density, (b) entanglement entropy and (c) magnetization in the same ensembles as Fig. 1. Data are binned in 100 bins. Parameters: $N=40$, $\Omega=2$, $\kappa t=8$.
Our results also highlight a crucial limitation of mean-field theory in capturing nonstabilizerness. While the mean-field description correctly predicts the behavior of magic in the magnetized phase, it fails in the BTC phase due to the mixed nature of the unconditional steady state. We have shown that quantum trajectories provide an alternative framework to describe magic in this regime, revealing its extensive scaling with system size and its independence from the unraveling scheme.
@@ Acknowledgements.
Appendix A Magic, entanglement and magnetization distributions across the unravellings
To gain further insights into the observation that the stabilizer entropy is unraveling-independent while the entanglement entropy is not, we also studied the magic distributions on the different ensembles of quantum trajectories shown in Fig. 1. For instance, in Fig. 9(a) we show the histogram of the density of magic of the $M=2000$ trajectories of the three ensembles, at $\kappa t=8$ (marked in the panels of Fig. 1 with vertical dotted lines). The histograms represent the probability density of a certain value of the magic density $m_{2}^{(k)}$ on the state |$\ket{\psi_{k}}$ of the unraveling considered. Besides having a similar mean, the distributions look identical. This is in line with the behavior of the magnetization, see Fig. 9(c), and completely different from the entanglement entropy, Fig. 9(b). These observations equally apply to other times $\kappa t$.

### v3

@@ VI Conclusions
At the critical point of the measurement-induced phase transition, the two complexity measures reveal complementary aspects of the behavior of the system. The entanglement entropy exhibits a genuine phase transition with unraveling-dependent scaling, while nonstabilizerness displays a cusp singularity in its derivative while, importantly, remaining extensive in system size across both phases. This result suggests that the superradiant model hosts highly nonclassical resources, even in the absence of entanglement.
Our results also highlight a crucial limitation of mean-field theory in capturing nonstabilizerness. While the mean-field description correctly predicts the behavior of nonstabilizerness in the magnetized phase, it fails in the BTC phase due to the mixed nature of the unconditional steady state. We have shown that quantum trajectories provide an alternative framework to describe nonstabilizerness in this regime, revealing its extensive scaling with system size and its independence from the unraveling scheme.
@@ Acknowledgements.
Appendix A Nonstabilizerness, entanglement and magnetization distributions across the unravellings
To gain further insights into the observation that the stabilizer entropy is unraveling-independent while the entanglement entropy is not, we also studied the stabilizer entropy distributions on the different ensembles of quantum trajectories shown in Fig. 1. For instance, in Fig. 9(a) we show the histogram of the density of SRE of the $M=2000$ trajectories of the three ensembles, at $\kappa t=8$ (marked in the panels of Fig. 1 with vertical dotted lines). The histograms represent the probability density of a certain value of the nonstabilizerness density $m_{2}^{(k)}$ on the state |$\ket{\psi_{k}}$ of the unraveling considered. Besides having a similar mean, the distributions look identical. This is in line with the behavior of the magnetization, see Fig. 9(c), and completely different from the entanglement entropy, Fig. 9(b). These observations equally apply to other times $\kappa t$.